### v2

# Modeling User Preferences as Distributions for Optimal Transport-based Cross-domain Recommendation under Non-overlapping Settings

Abstract
Cross-domain recommender (CDR) systems aim to transfer knowledge from data-rich domains to data-sparse ones, alleviating sparsity and cold-start issues present in conventional single-domain recommenders. However, many CDR approaches rely on overlapping users or items to establish explicit cross-domain connections, which is unrealistic in practice. Moreover, most methods represent user preferences as fixed discrete vectors, limiting their ability to capture the fine-grained and multi-aspect nature of user interests.
To address these limitations, we propose DUP-OT (Distributional User Preferences with Optimal Transport), a novel framework for non-overlapping CDR. DUP-OT consists of three stages: (1) a shared preprocessing module that extracts review-based embeddings using a unified sentence encoder and autoencoder; (2) a user preference modeling module that represents each user’s interests as a Gaussian Mixture Model (GMM) over item embeddings; and (3) an optimal-transport-based alignment module that matches Gaussian components across domains, enabling effective preference transfer for target-domain rating prediction.
Experiments on Amazon Review datasets demonstrate that DUP-OT mitigates domain discrepancy and significantly outperforms state-of-the-art baselines under strictly non-overlapping training settings, with user correspondence revealed only for inference-time evaluation. Code is available at https://github.com/XiaoZY2000/DUP-OT-organized.
Keywords Cross-Domain Recommendation $\cdot$ Non-overlapping Users $\cdot$ Gaussian Mixture Model (GMM) $\cdot$ Optimal Transport $\cdot$ Cold-start Problem
1 Introduction
With the rapid growth of online information, recommender systems have become essential for helping users navigate massive content R_Lü_2012. However, conventional single-domain recommenders suffer from cold-start and data sparsity issues T_He_2024; L_Liu_2024; A_Zang_2023. Cross-domain recommendation (CDR) alleviates these problems by transferring knowledge from a dense source domain to a sparse target domain A_Zhu_2018. Yet most existing CDR methods rely on overlapping users or items in the training stage to bridge domains and typically represent user preferences as discrete vectors, which limits their ability to capture multi-aspect preferences.
In this work, we study the non-overlapping CDR scenario—where no users or items are shared across domains—and propose to represent user preferences using Gaussian Mixture Models (GMMs). This enables fine-grained, multi-aspect preference modeling, but also introduces two challenges: (1) domain discrepancy becomes harder to address without overlapping entities D_Guo_2024; C_Liu_2023, and (2) GMM-based preference modeling significantly increases computation and is not directly compatible with standard optimal transport formulations.
Recent CDR approaches attempt to reduce domain discrepancy using contrastive learning, adversarial training, or distribution alignment L_Liu_2024; R_Wang_2020; C_Perera_2019. Other studies explore distributional representations but either adopt simple Gaussian models or still rely on overlapping entities M_Zhu_2024. Thus, how to build expressive GMM-based preference representations while enabling cross-domain alignment in a strictly non-overlapping setting remains an open problem.
To address this, we propose DUP-OT (Distributional User Preferences with Optimal Transport), a novel framework that models each user’s preference as a GMM over item representations and aligns domains using optimal transport (OT). To reduce computational cost and make OT feasible, we assume that all users within a domain share a domain-level set of Gaussian components extracted from item representations, while each user learns personalized weights over these components. DUP-OT consists of three stages: (1) a shared preprocessing stage with a unified sentence encoder and autoencoder to obtain low-dimensional item and user embeddings; (2) a user preference modeling stage where each domain fits item-level GMM components and learns user-specific mixture weights; and (3) an OT-based alignment stage that transports GMM weights from the source to the target domain to enhance target-domain rating prediction.
Our contributions are summarized as follows:
  • We highlight the limitations of discrete vector representations in non-overlapping CDR and motivate the need for distributional preference modeling.
  • We propose DUP-OT, a GMM-based framework that aligns cross-domain preferences using optimal transport on distributional representations.
  • We validate DUP-OT on Amazon Review datasets, showing consistent performance gains over baselines.
2 Related Works
Cross-Domain Recommendation. CDR aims to transfer knowledge from a rich source domain to a sparse target domain. Early approaches rely on overlapped users or items through representation alignment D_Li_2020, adversarial learning R_Li_2022, and graph-based propagation D_Guo_2021. However, such overlaps are rarely available in practice.
Non-overlapping CDR methods typically learn shared latent spaces using embedding alignment L_Wang_2021, generative modeling T_Salah_2021, or distribution matching M_Zhu_2024. While recent work explores Gaussian-based representations, they either use simple single-Gaussian models or still rely on partial overlaps. Modeling user preferences with expressive GMMs under fully non-overlapping settings remains underexplored.
Optimal Transport in CDR. OT provides a principled tool for aligning distributions across domains. It has been applied to match user/item embeddings G_Li_2022 or latent distributions L_Liu_2024. However, prior work does not integrate OT with GMM-based user preference modeling, nor address the computational challenges in fully non-overlapping scenarios.
3 Preliminary
Definitions of notations. We consider two domains: the source domain $s$ and the target domain $t$. For each domain $x\in\{s,t\}$, we denote the user set and item set as $\mathcal{U}_{x}$ and $\mathcal{V}_{x}$. Although users or items may exist in both domains in reality, in our CDR setting we impose that the model does not use any overlapping-user or overlapping-item information during training. Therefore, the training process treats the two domains as non-overlapping, even if overlaps may exist in the raw data.
For the $i$-th user and $j$-th item in domain $x$, each interaction is written as:
$(u_{i}^{x},\,v_{j}^{x},\,r_{ij}^{x},\,h_{ij}^{x},\,t_{ij}^{x}),$
where $r_{ij}^{x}$ is the rating, $h_{ij}^{x}$ is the corresponding review text and $t_{ij}^{x}$ is the time stamp.
Target-domain dataset split. The target-domain interactions are divided into a training set $\mathcal{D}_{t}^{\mathrm{train}}$, a validation set $\mathcal{D}_{t}^{\mathrm{valid}}$, and a test set $\mathcal{D}_{t}^{\mathrm{test}}$. During training, the model only uses $\mathcal{D}_{s}^{\mathrm{train}}$ and $\mathcal{D}_{t}^{\mathrm{train}}$, and any user/item overlaps between the two domains are explicitly ignored. However, validation and test sets may contain users or items that also appear in the source domain, since they are used only for evaluation and not for model optimization.
Practical motivation. This setting corresponds to real-world scenarios where user or item identities might be shared across platforms, but privacy constraints or system limitations prevent such cross-domain links from being used during training. For example, users may later appear in the target domain after the model has been trained, or items may be synchronized across platforms only at deployment time. Thus, overlaps can exist for evaluation but are not exploited for learning, faithfully reflecting practical cross-domain recommendation conditions.
Task Formulation. We aim to address the cold-start problem in the target domain by transferring knowledge from the source domain under a non-overlapping CDR setting. Specifically, the goal is to leverage source-domain data to enhance the rating prediction performance in the target domain.
4 Methodology
4.1 Overview
Figure 1: The overall structure of DUP-OT
Figure 1 illustrates the architecture of our proposed DUP-OT framework, which enables non-overlapping cross-domain recommendation based on distributional user preference representations. The pipeline consists of three main stages: (1) shared preprocessing, (2) user GMM weight learning, and (3) cross-domain rating prediction.
In the shared preprocessing stage, we extract and align user and item embeddings across domains. Specifically, we employ a shared pre-trained sentence encoder to encode reviews into initial user and item embeddings, followed by a shared autoencoder trained on both domains to reduce embedding dimensionality and ensure a unified feature space.
The user GMM weight learning stage is performed separately in each domain. We first fit a Gaussian Mixture Model (GMM) on the item embeddings to obtain shared Gaussian components. Then, an MLP is trained to map each user embedding to a set of weights over these components, forming a user-specific GMM. Another MLP is trained to predict user-item ratings based on the Mahalanobis distance between the item embedding and each Gaussian component weighted by the user’s distribution.
Finally, in the cross-domain rating prediction stage, we align the GMM components across domains via Optimal Transport (OT) and use the resulting transport matrix to transfer users’ GMM weights from the source domain to the target domain. For users in the target-domain test set, our framework handles three cases: if a user has interactions in both domains during training, we combine the transferred distribution with the target-domain distribution; if the user has interactions only in the source domain, we use the transferred distribution; and if the user has interactions only in the target domain, we use the target-domain distribution. These resulting user distributions are then used to enhance rating prediction performance in the target domain.
4.2 Shared Preprocessing Stage
To capture semantic similarity across domains, we derive user and item embeddings from review text. For each user and item, all associated reviews are encoded using a shared pre-trained sentence encoder, and we apply a time-aware weighting scheme so that recent reviews contribute more to the final representation.
The encoder outputs high-dimensional vectors, which make GMM fitting unstable. Therefore, we train a shared autoencoder on both domains to obtain compact, domain-consistent embeddings. This unified latent space preserves semantic structure while reducing dimensionality, providing stable inputs for subsequent GMM construction and cross-domain alignment.
4.3 User GMM Weights Learning Stage
Users often exhibit multi-aspect preferences, which discrete vector representations struggle to capture. Therefore, DUP-OT models each user’s preference as a Gaussian Mixture Model (GMM) over item embeddings.
To reduce computational cost, we assume that all users within a domain share a domain-level set of Gaussian components. These components are obtained by fitting a GMM on item embeddings using the EM algorithm (implemented via BayesianGaussianMixture). Each user is then mapped to a personalized set of mixture weights over these components.
For each domain, we train two MLPs: a weight learner that predicts user-specific mixture weights from user embeddings, and a rating predictor that takes the weighted component-wise Mahalanobis distances between a user and an item to estimate the final rating. The learning objective is the standard mean squared error:
$\mathbf{w}^{x}_{u_{i}}$ $=\text{W-Learner}^{x}(\mathbf{z}^{x}_{u_{i}}),$ (1)
$\hat{r}^{x}_{ij}$ $=\text{R-Predictor}^{x}\big(\mathbf{w}^{x}_{u_{i}},\,\mathbf{z}^{x}_{v_{j}},\,\mu^{x},\,\Sigma^{x}\big),$ (2)
$\mathcal{L}_{\mathrm{rating}}^{x}$ $=\frac{1}{|\mathcal{I}^{x}|}\sum_{(i,j)\in\mathcal{I}^{x}}\!\!\left(r_{ij}-\hat{r}_{ij}^{x}\right)^{2}.$ (3)
Here, $\mu^{x}$ and $\Sigma^{x}$ denote the domain-level Gaussian component means and covariances. This stage produces a distributional user representation that is later aligned across domains via optimal transport.
4.4 Cross Domain Rating Prediction Stage
In this stage, we enhance the target-domain rating predictor by constructing improved user distributions through Optimal Transport (OT). This stage consists of two parts: aligning the GMM components across domains using OT, and generating enhanced user-specific distributions in the target domain according to three test-time scenarios.
OT-based alignment of Gaussian components. Since all users within a domain share the same set of Gaussian components, we apply OT only between the two domains’ GMM component sets, avoiding instance-level transport. The OT problem is formulated as Eq. 4.
$\text{arg}\min_{\mathcal{T}}$ $\sum_{i=1}^{m}\sum_{j=1}^{n}\mathcal{T}_{ij}\mathcal{C}_{ij}$ (4)
s.t. $\sum_{i=1}^{m}\mathcal{T}_{ij}=\frac{1}{n},\quad\sum_{j=1}^{n}\mathcal{T}_{ij}=\frac{1}{m}$
where $m$ and $n$ denote the numbers of Gaussian components in the source and target domains. The cost matrix $\mathcal{C}$ is computed using the Wasserstein-2 distance (Eq. 5), and we solve for the optimal transport matrix $\mathcal{T}$ using the Sinkhorn algorithm.
$\mathcal{W}_{2}^{2}(\mathcal{N}_{1},\mathcal{N}_{2})=\;$ $\left\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\right\|_{2}^{2}$
$+\operatorname{Tr}\left(\boldsymbol{\Sigma}_{1}+\boldsymbol{\Sigma}_{2}-2\left(\boldsymbol{\Sigma}_{2}^{1/2}\boldsymbol{\Sigma}_{1}\boldsymbol{\Sigma}_{2}^{1/2}\right)^{1/2}\right)$ (5)
User weights are then transported as Eq. 6.
$\mathbf{w}^{t}_{u_{i}}=\mathbf{w}^{s}_{u_{i}}\mathcal{T}$ (6)
Three test-time cases. For each user in the target-domain test set, we construct a final user distribution in the target domain depending on the interaction availability:
  • Both domains have interactions: we linearly fuse the transferred distribution and the target-domain distribution to form the final user distribution:
    $\mathbf{w}_{u_{i}}^{\text{final}}=\alpha\,\mathbf{w}_{u_{i}}^{t}+(1-\alpha)\,\mathbf{w}_{u_{i}}^{t,\text{orig}},$
    where $\alpha\in[0,1]$ controls the contribution of the transferred distribution.
  • Only source domain has interactions: use the transferred distribution alone:
    $\mathbf{w}_{u_{i}}^{\text{final}}=\mathbf{w}_{u_{i}}^{t}.$
  • Only target domain has interactions: use the target-domain distribution alone:
    $\mathbf{w}_{u_{i}}^{\text{final}}=\mathbf{w}_{u_{i}}^{t,\text{orig}}.$
For users without interactions in either domain, we initialize their mixture weights with random values. The resulting user distribution is then fed into the target-domain rating predictor to produce the final enhanced rating prediction.
5 Experiments and Results
In this section, we conduct experiments on our proposed model and several baseline models and compare the results to validate our model’s effectiveness. Specifically, the research questions we focus on are as follows:
RQ1. Does introducing cross-domain information enhance the recommendation performance in the target domain?
RQ2. Does modeling user preferences as distributions, rather than vectors, lead to improved recommendation performance?
RQ2. How does the performance of our proposed method compare with that of existing CDR models?
To answer RQ1, we conduct an ablation study comparing target-domain performance with and without source-domain information. To address RQ2, we evaluate our target-domain model (using distribution-based user preferences but without source-domain data) against single-domain baselines. To address RQ3, we compare our proposed DUP-OT model with representative cross-domain recommendation baselines.
5.1 Experimental Setup
In this section, we introduce some details about our experiments’ setup, including datasets, model setting, baseline models, metrics.
Table 1: Results of baseline models and DUP-OT
(a) Results on the Domain Pair $(D_{s},D_{t})$
$D_{s}\to D_{t}$ Digital Music $\to$ Electronics Movies & TV $\to$ Electronics Video Games $\to$ Electronics
Model RMSE MAE RMSE MAE RMSE MAE
TDAR 1.5688 1.0904 1.5677 1.0896 1.5688 1.0902
DUP-OT (w/o source) 1.3699 0.9463 1.3494 0.8806 1.3716 0.9481
DUP-OT (w/ source) 1.2919 0.8965 1.2907 0.8774 1.3376 1.0032
(b) Single-domain recommendation results (Electronics-only)
Model RMSE MAE
LightGCN 1.5317 1.1179
NeuMF 1.4599 1.3297
DUP-OT (w/o source average) 1.3636 0.9250
5.1.1 Datasets
We use the Amazon Review 5-core datasets J_Ni_2019: Digital Music, Movies and TV, Video Games, and Electronics. Electronics is used as the target domain, and the other three as source domains. Target-domain data are split into train/validation/test (8:1:1) chronologically. To keep temporal consistency, we filter out source-domain interactions occurring later than the earliest validation timestamp in the target domain. The filtered source data and target-domain training set are used to train DUP-OT, and evaluation is performed on the target-domain test set.
5.1.2 Model Setting
Our proposed DUP-OT consists of three stages. In the first stage, we use a shared pre-trained sentence encoder (all-MiniLM-L6-v2) from Sentence-Transformers to obtain review-based user and item embeddings, followed by a shared autoencoder to reduce dimensionality and form a unified latent space across domains. In the second stage, for each domain we fit item-level Gaussian Mixture Models (GMMs) and train two small MLPs: one to predict user-specific mixture weights over the shared Gaussian components, and one to predict ratings based on the resulting user distributions and item embeddings. The number of GMM components $K$ is determined automatically by the BayesianGaussianMixture estimator.
5.1.3 Training Settings
We train all models using the Adam optimizer with a learning rate of $3\times 10^{-4}$ and a batch size of 256, with early stopping based on target-domain validation loss. The sentence encoder is frozen, while the shared autoencoder is jointly trained across domains to obtain unified low-dimensional embeddings. For each domain, we fit item-level Gaussian Mixture Models using the BayesianGaussianMixture estimator and train domain-specific MLPs for predicting user GMM weights and ratings. Users appearing only in validation or test sets are treated as cold-start users. All models are implemented in PyTorch and trained on a single NVIDIA GPU. To ensure robustness, all experiments are run with five random seeds $\{7,11,17,33,2025\}$, and we report the mean of all metrics.
5.1.4 Baselines
We implement the following models as baseline models for comparison:
Single Domain Models:
[…]
Non-overlapping Cross Domain Models:
[…]
For single domain RS models, we utilize RecBole recbole[1.0]; recbole[2.0]; recbole[1.2.1] in our experiments to implement them. And for TDAR, we just manually implement them.
5.1.5 Ablation Study: Effect of Source-Domain Information
To evaluate whether source-domain information enhances recommendation performance in the target domain (RQ1), we compare two variants of our framework:
  • DUP-OT (w/ source): the full model that uses both domains. The autoencoder is trained jointly on both domains, and in the second stage the source-domain GMMs, GMM-weight learners, and rating predictors are trained using source-domain data. User distributions are then transferred to the target domain via Optimal Transport and fused with target-domain distributions during inference.
  • DUP-OT (w/o source): an ablated model that does not use any source-domain interactions during the second stage. The autoencoder is still trained jointly on both domains to ensure consistent representation spaces, but the GMM construction, GMM-weight learner, and rating predictor are trained exclusively on the target-domain training set. No cross-domain alignment or distribution transfer is performed.
This setup ensures that the only difference between the two variants lies in whether cross-domain preference information is introduced.
5.1.6 Metrics
To evaluate the performance of our proposed DUP-OT and baseline models, we set all models’ targets as rating prediction on the target domain’s test dataset. And then we use RMSE and RAE as metrics to evaluate models’ performance. The results are shown in Table 1.
5.2 Discussion of Results
Based on the results in Table 1, we now answer the three research questions raised at the beginning of this section.
[…]
Answer to RQ2 (Effect of distribution-based user modeling). Comparing DUP-OT(w/o source) with single-domain recommenders (LightGCN and NeuMF), DUP-OT(w/o source) achieves substantially better performance in the Electronics-only setting, despite using no source-domain data. This indicates that modeling user preferences as distributions rather than discrete vectors yields more expressive representations and improves rating prediction accuracy in the target domain.
Answer to RQ3 (Comparison with cross-domain baselines). Relative to TDAR, which is a representative CDR method, DUP-OT(w source) achieves consistently lower RMSE and MAE across all dataset combinations. This demonstrates that combining GMM-based preference modeling with optimal transport provides more effective cross-domain alignment than conventional approaches, validating the overall superiority of our framework.
6 Conclusion
We proposed DUP-OT, a distribution-based framework for non-overlapping cross-domain recommendation. By modeling user preferences as GMMs and aligning domains via Optimal Transport, DUP-OT enables expressive preference representation and effective cross-domain transfer without relying on shared users or items during the training stage. Experiments on Amazon Review datasets show that DUP-OT improves target-domain rating prediction, outperforms single-domain models even without source data, and achieves superior performance over a representative non-overlapping CDR baseline.